OpenStudy (anonymous):

A 2.00 g bullet is fired from a rifle with an in itial velocity of 841 m/s. If the rifle has a mass of 3.80 kg, the recoil speed of the rifle

OpenStudy (anonymous):

i did 0=mv+m2v2

OpenStudy (anonymous):

i was getting 4.42*10^-2 but thats wrong ?

OpenStudy (anonymous):

i get 4,42*10^-2 too

OpenStudy (anonymous):

the answer is 3144

OpenStudy (anonymous):

like 3.14*10^-4

OpenStudy (anonymous):

@IrishBoy123 what u thiink?

OpenStudy (irishboy123):

your method seems sound though you are out by a factor of 10. \(4.42(10^{-2}) \times 3.8 = 0.168\) \(2(10^{-3}) \times 841 = 1.68\) i just wonder if it's worth posting a scan of the question or a link...

OpenStudy (anonymous):

@IrishBoy123 here

OpenStudy (irishboy123):

so that gives \(4.43 \times 10^{-1}\) or in the form the question requires: a 4 b 4 c 3 d 1 which starts to *resemble* the answer you posted, when you posted it as "3144"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the answer is wrong ?

OpenStudy (anonymous):

like a typo u think?

OpenStudy (anonymous):

@IrishBoy123 can u help me with this ?

OpenStudy (irishboy123):

sure post as a new question if you can do that now, i can be here